### v1

@@ 4 Conclusions
We have presented new techniques to solve nonlinear, multiscale differential equations (DEs) with Physics-Informed Neural Networks (PINNs), placing special emphasis on inverse problems. These techniques, which we term multi-head (MH) and Unimodular Regularization (UR), have proved effective in various examples of nonlinear, multiscale DEs. We have shown that, by using a single NN structure (the body) shared across multiple NN substructures (heads), MH training can learn the broader space of solutions for variations of the DE. We verified this approach using variations in the DEs’ parametric dependence, initial conditions (ICs), and boundary conditions (BCs). The learned general solution space (or latent space) can then be leveraged to find specific solutions for new DE variations under more extreme or stiffer regimes, even if the MH body was not initially trained for these cases. This approach is known as transfer learning (TL). Furthermore, we have developed UR, a technique that imposes specific conditions on the determinant of the metric tensor induced by embedding the latent space into the input/parameter space. This idea is grounded in differential geometry and draw inspiration from General Relativity. The imposed conditions quantify and regulate the latent space’s sensitivity to specific variations in the DEs. We have found that combining the MH approach with UR in the latent space of solutions significantly enhances PINN efficiency, facilitating the TL process and thus enabling more effective solution discovery.
@@
Our results are derived by implementing the MH technique both with and without UR, then comparing the solutions obtained to a numerical baseline. We apply this approach to three distinct sets of ordinary differential equations (ODEs): the flame equation, the van der Pol (damped) oscillator, and the Einstein Field Equations (EFEs).
@@
For the flame equation, a nonlinear first-order ODE, UR reduces the root mean square error (rms) from 35.1% to 0.5%. For the van der Pol oscillator, a nonlinear second-order ODE, UR enables transfer learning to highly stiff regimes, resulting in a root mean square error (rms) of 5.7%. Finally, for the Einstein Field Equations (EFE)—an inverse, nonlinear system of first-order ODEs—we can simultaneously recover an approximate free-form function that appears in the DEs, together with the corresponding solutions. The free function is inferred from its physically motivated correspondence to a physical relation between different BCs (an implicit mapping in the equations). In this final case, we applied transfer learning to two increasingly stiff regimes. The first transfer showed a factor-of-2 improvement with UR. The second demonstrated a factor-of-3 improvement—or factor-of-7 when comparing the recovered free-form function or the corresponding BC relation. In this second instance, only approximate recovery of the full solution’s morphology is achieved when UR is included.
@@
We have provided three examples demonstrating that, with these algorithms, it is possible to tackle very complex problems in various scientific fields where unknown phenomena must be discovered through the inversion of differential equations, as shown in the last provided example. Consequently, these algorithms could be extremely powerful, as such problems arise in a wide range of areas, including General Relativity, complex fluid dynamics, social sciences, and beyond.

### v2

@@ 4 Conclusions
We have presented two techniques to solve nonlinear, multiscale differential equations (DEs) with Physics-Informed Neural Networks (PINNs), placing special emphasis on inverse problems. These techniques, which we term multi-head (MH) and Unimodular Regularization (UR), have proved effective in various examples of nonlinear, multiscale DEs. We have shown that, by using a single NN structure (the body) shared across multiple NN substructures (heads), MH training can learn the broader space of solutions for variations of the DE. We verified this approach using variations in the DEs’ parametric dependence, initial conditions (ICs), and boundary conditions (BCs). The learned general solution space (or latent space) can then be leveraged to find specific solutions for new DE variations under more extreme or stiffer regimes, even if the MH body was not initially trained for these cases. This approach is known as transfer learning (TL). Furthermore, we have developed Unimodular Regularization (UR), a technique that imposes specific conditions on the determinant of the metric tensor induced by embedding the latent space into the input/parameter space. This idea is grounded in differential geometry and draw inspiration from General Relativity. The imposed conditions quantify and regulate the latent space’s sensitivity to specific variations in the DEs. We have found that combining the MH approach with UR in the latent space of solutions significantly enhances PINN efficiency, facilitating the TL process and thus enabling more effective solution discovery.
@@
Our results are derived by implementing the MH technique both with and without UR, then comparing the solutions obtained to a numerical baseline. We apply this approach to three distinct sets of ordinary differential equations (ODEs): the flame equation, the van der Pol (damped) oscillator, and the Einstein Field Equations (EFEs), being the last an inverse problem.
@@
For the flame equation, a nonlinear first-order ODE, UR reduces the root mean square error (RMS) from 35.1% to 0.5%. For the van der Pol oscillator, a nonlinear second-order ODE, UR enables TL to highly stiff regimes, resulting in a reduction of the RMS from 52.4% to 5.7%. Finally, for the Einstein Field Equations (EFE)—an inverse, nonlinear system of first-order ODEs—we can simultaneously recover an approximate free-form function that appears in the DEs, together with the corresponding solutions. The free function is inferred from its physically motivated correspondence to a physical relation between different BCs (an implicit mapping in the equations). In this final case, we applied TL to two increasingly stiff regimes. The first transfer showed a factor-of-2 improvement with UR. The second demonstrated a factor-of-3 improvement—or factor-of-7 when comparing the recovered free-form function or the corresponding BC relation. In this second instance, only approximate recovery of the full solution’s morphology is achieved when UR is included.
Although we have only considered ODEs, the extension of our setup to Partial Differential equations (PDEs) is almost straightforward. In the presented setup, promoting the system from ODEs to PDEs by adding a new independent variable is the same as considering additional Bundle parameters, since in both cases these are just inputs to the model. However, this convenience comes at a cost: for higher-dimensional PDEs or bundle solutions, we must sample the expanded input space on a multidimensional grid, which raises the effective batch size and correspondingly increases training time.
Furthermore, applying the UR method has some disadvantages and computational limitations. The first one is the training time. By including UR one needs to compute additional gradients during training in order to compute the metric tensor of the latent space. This increases the training time by $5\%-10\%$, but this number is highly dependent on $N$ and the batch size. Second, storing the metric tensor and its derivatives for every batch element can exhaust GPU memory (in our final EFE experiment, adding one extra head exceeded the capacity of an NVIDIA A30 GPU). To address these issues, future work will explore low‐rank or sparse tensor approximations, gradient checkpointing or selective gradient retention, and optimized CUDA kernels to reduce both memory usage and runtime.
@@
We have provided three examples demonstrating that, with these algorithms, it is possible to tackle very complex problems in various scientific fields where unknown phenomena must be discovered through the inversion of differential equations, as shown in the last provided example. In the future, we expect to apply this method to two different situations. The first one is to try to recover the potential in the holography setup in the most stiff cases. This will allow us to access physical situations with a very rich phenomenology. The second one is to apply the MH setup to gain some insights about the phenomena of turbulence. However, as we have explained, going directly to the Navier-Stokes equation is both computationally and conceptually demanding. It is for this reason that we are currently applying the set-up first to simpler PDEs, such as Burger’s equation. This will allow us to recover the basis elements that the NN is choosing to solve the equation, up to a certain precision. More generally, we consider these algorithms could be extremely powerful and generalizable to a wide range of areas, including General Relativity, complex fluid dynamics, social sciences, and beyond.
5 Author contribution
All the authors have contributed equally to this work. Specifically all the authors have actively participated on the conceptual idea, writing code, training models and writing the final manuscript.
6 Data Availability
The data points used for BCs in subsection “Einstein Field Equations” are available in the GitHub repository https://github.com/pedrota2000/Unimodular_regularization. The trained models are available upon request to the authors.
7 Code availability
The Python codes used to generate the results are available in the GitHub repository https://github.com/pedrota2000/Unimodular_regularization.
8 Competing Interests
The authors declare no competing interests.
@@ Acknowledgements
Supplementary Note 1: Mathematical form of the metric for UR and comparison to Jacobian Regularization
In this section, we mathematically show that a Jacobian regularization [^1] Note that we changed notation in the choice of index in $J_{c,i}(\mathbf{x})$ from [38], to $J_{i,\mu}(\mathbf{x})$ to make contact with the index notation in the subsection “Unimodular Regularization” within the “Methods” section, and throughout this work. (JR) of the Frobenius norm of the pseudo-Jacobian [38]:
$L_{\text{JR}}=\lambda_{\text{JR}}\sum_{\text{batch}}||J(x)||^{2}_{F}=\lambda_{\text{JR}}\sum_{\text{batch}}\sum_{i=1}^{d}\sum_{\mu=1}^{n}\left[J_{i,\mu}(\mathbf{x})\right]^{2}\,\,\,\,\,\,;\,\,\,\,\,J_{i,\mu}(\mathbf{x})\equiv\left.\frac{\partial H_{i}}{\partial x^{\mu}}\right|_{\mathbf{x}}$ (16)
is implicit in UR, with $n$ inputs and a dimension of the latent space $d$ (number of “categories” in [38]). This is because it appears as one of the terms in the computation of the determinant of the metric that is then regularized to unity, i.e.:
Before the computation, we make the following remark: the Frobenious norm of the pseudo-Jacobian matrix is defined such that all the partial derivatives with respect to inputs $x^{\mu}$ with $\mu=1,\dots,n$ are minimized individually. However, this scalar quantity lacks an interpretation in geometrical terms. In fact, we call the matrix $J_{i,\mu}\left(\textbf{x}\right)$ a pseudo-Jacobian because it is not generally a square matrix, and therefore cannot be the geometrical Jacobian as is usually understood in the literature. On the other hand, the minimization of the determinant of the metric, $g=\text{det}(g_{\mu\nu})\rightarrow 1$ (through UR) has a clear geometrical interpretation; the determinant of the metric, or geometric Jacobian, measures the volume of a differential of hyper-volume in the latent space $H_{i}(x^{\mu})$, where $i=1,\dots,d$, being $d$ the dimension of the latent space. Regularizing this quantity to unity leads to transforming the latent space such that this differential of hyper-volume is smoothly responsive to all inputs $x^{\mu}$, while keeping this response small if the change in $x^{\mu}$ is also small. This second point is achieved by minimizing the derivatives of the latent space with respect to all the inputs (similar to Jacobian regularization, but in a way that preserves the geometrical interpretation).
To show this, let us explicitly compute $g=\text{det}(g_{\mu\nu})$. The metric in our geometric construction is defined as:
$g_{\mu\nu}\,=\,\frac{\partial\vec{\Omega}}{\partial x^{\mu}}\cdot\frac{\partial\vec{\Omega}}{\partial x^{\nu}}$ (18)
being
$\vec{\Omega}\,=\,\left(x^{\mu},H_{i}(x^{\mu})\right)=(x^{1},\dots,x^{n},H_{1}(x^{\mu}),\dots,H_{d}(x^{\mu}))\vskip 2.84544pt$
where $x^{\mu}\in\mathbb{R}^{n}$ are the inputs to the NN (independent variable, IC/BC, parameters), and $H\in\mathbb{R}^{d\times n}$ is the latent space given by the output of the last hidden layer of the body of the NN. This metric is computed for every point in the batch.
We denote the derivatives with respect to all inputs as:
$\frac{\partial}{\partial x^{\mu}}\equiv\left(\frac{\partial}{\partial x^{1}},\ldots,\frac{\partial}{\partial x^{n}}\right)\equiv(\partial_{1},\ldots,\partial_{n})$
Then:
$\frac{\partial\bar{\Omega}}{\partial x^{\mu}}=\begin{pmatrix}\partial_{1}\\ \vdots\\ \partial_{n}\end{pmatrix}\cdot\left(x^{1},\ldots,x^{n},H_{1}(x^{\mu}),\ldots,H_{d}(x^{\mu})\right)=\begin{pmatrix}1&\cdots&0&\partial_{1}H_{1}&\cdots&\partial_{1}H_{d}\\ \vdots&\ddots&\vdots&\vdots&\ddots&\vdots\\ 0&\cdots&1&\partial_{n}H_{1}&\cdots&\partial_{n}H_{d}\end{pmatrix}_{n\times(n+d)}$
Computing then the metric in matrix form:
$\left(g_{\mu\nu}\right)_{n\times n}=\left(\frac{\partial\vec{\Omega}}{\partial x^{\mu}}\right)\cdot\left(\frac{\partial\vec{\Omega}}{\partial x^{\nu}}\right)^{T}=\begin{pmatrix}1+\sum\limits_{i=1}^{d}a_{1i}^{2}&\cdots&\sum\limits_{i=1}^{d}a_{ni}a_{1i}\\ \vdots&\ddots&\vdots\\ \sum\limits_{i=1}^{d}a_{1i}a_{ni}&\cdots&1+\sum\limits_{i=1}^{d}a_{ni}^{2}\end{pmatrix}$ (19)
where we defined $a_{\mu i}\equiv\partial_{\mu}H_{i}$. This result can also be written as:
$g_{\mu\nu}=\left\{\begin{array}[]{lcl}1+\sum_{i=1}^{d}\left(\partial_{\mu}H_{i}\right)^{2}&;&\mu=\nu\\ &&\\ \sum_{i=1}^{d}\left(\partial_{\mu}H_{i}\right)\left(\partial_{\nu}H_{i}\right)&;&\mu\neq\nu\end{array}\right$. (20)
The expression of the determinant for general $n,d$ becomes lengthy. For the purpose of this work, let us show an example for the case n=2,d=2n=2,\,d=2, where we have the determinant:
$g=\text{det}\left(g_{\mu\nu}^{(2,2)}\right)$ $=1+\left(\sum_{\mu=1}^{n}\sum_{i=1}^{d}a_{\mu i}^{2}\right)+\left(a_{11}a_{22}-a_{12}a_{21}\right)^{2}$
$=1+\sum_{\mu=1}^{n}\sum_{i=1}^{d}\left[J_{i,\mu}(\mathbf{x})\right]^{2}+\left(a_{11}a_{22}-a_{12}a_{21}\right)^{2}=$
$\equiv 1+J^{2}+A^{2}$ (21)
where $A^{2}=\left(a_{11}a_{22}-a_{12}a_{21}\right)^{2}>0$. In the second line we used the definition of the matrix terms of the pseudo-Jacobian matrix used in [38], $a_{\mu i}=J_{i,\mu}(\mathbf{x})\equiv\partial H_{i}/\partial x^{\mu}$. Thus, we can see that regularizing $g\rightarrow 1$ (we actually do $\sqrt{g}\rightarrow 1$) is equivalent to minimizing the factor $(J^{2}+A^{2})$, which is a regularization that includes JR (minimizing the Frobenius norm $J^{2}$), together with an extra regularizing factor.
This calculation can be generalized to arbitrary number of inputs $n$ and dimension of the latent space $d$.
Supplementary Note 2: Relation of UR with the Lipschitz constant.
The Lipshitz constant $L$ of a function $f$ is defined as
$\frac{|f(y)-f(x)|}{|y-x|}\leq L$
When considering input to solution mapping $u(x^{\mu})$ we can write the this condition as follows
$\frac{|u(y^{\mu})-u(x^{\mu})|}{|y^{\mu}-x^{\mu}|}\leq L$
It is easy to prove that this quantity is bounded by the 2-norm of the Jacobian of the function $u$
[…]
Note that by using the chain rule we can relate the derivative of $u$ with respect to the inputs to the derivative of the latent space as
$\frac{\partial u(x^{\mu})}{\partial x^{\nu}}\,=\,\sum_{i}\frac{\partial u(x^{\mu})}{\partial H_{i}(x^{\mu})}\cdot\frac{\partial H_{i}(x^{\mu})}{\partial x^{\mu}}$
since the determinant of $g$ includes the derivatives of the latent space with respect to the inputs (as shown in “Supplementary Note 1”) making this quantity small would make the Lipschitz constant small.
Intuitively, we can think of this regularization techniques as follows. The determinant of the metric $g$ is proportional to the derivatives of the latent space with respect to the inputs. These derivatives are quantifying the sensitivity of the latent space to changes of independent variables and parameters of the differential equations. Let us consider two different limits: if the derivatives are very small, then the response of the latent space will be constant, and we will be learning no structure at all to project down to the actual solution. On the other hand, if the response of the latent space is very sensitive to the inputs, there will be too much structure in the latent space and it will be very difficult for the head to predict outside the range of training during the transfer learning. Thus, we will try to find the perfect balance between these two limits by playing with the hyperparameters $\lambda$ and $N$ (defined in the main text in the “Unimodular Regularization” subsection within the “Methods” section).